1. Entertainer Madonna weighs 52 kg and is 160 cm in height. Estimate her BSA.

2. Order: phenobarbitol sodium 400 mg po in two divided doses daily. The recommended dosage of this anticonvulsant is 1–3 mg/kg/d. Is the prescribed dose safe for the client who weighs 203 pounds?

3. Estimate the body surface area of a client who is 5 feet tall and weighs 60 kg.

4. Order: Percocet (oxycodone/acetaminophen) 10/650 mg po prn pain q6h. How many 5/325 mg tablets of this opioid analgesic will you administer to the client?
Answer #1

ANSWER.

1) Calculation of BSA.

          Formula

               BSA = Squareroot (height(cm) * Weight(kg)) / 3600)

     Here

          Patient weight = 52 kg.

                      Height = 160 cm

As per above formula

            BSA = Squareroot (160 * 52) / 3600)

                    = Squareroot (8320 / 3600)

                    = Squareroot (2.3)

                    = 1.5 m2.

2) Calculation of safe dose.

           Ordered dose = 400 mg.

           Recommended dose = 1-3 mg/kg/day.

          Weight of the child = 203 pounds.

Convert pound to kg for calculation

        1 pound = 0.454 kg

   Weight of the patient = 203 * 0.454

                                      = 92.16 kg.

     Dose required for this patient = 1 * 92.16 to 3 * 92.16

                                                     = 92.16 mg to 276 .5 mg per day.

So

       The ordered dose(400mg/day) is NOT SAFE for this patient.

3) Calculation of BSA.

        Formula

            BSA = Squareroot (Height(cm) * Weight(kg)) / 3600)

     Here

           Height = 5 feet.

    Convert 5 feet to cm for calculation

              1 feet = 30.5 cm

     So

          Height = 5 * 30.5

                     = 152.5 cm.

          Weight = 60 kg.

As per above formula

          BSA = Squareroot (152.5 * 60 / 3600).

                   = Squareroot (9150 / 3600)

                   = Squareroot (2.5)

                   = 1.58

                   = 1.6 m2.

4) Calculation of number of tablets required.

            Ordered dose = 10/650 mg

            Available dose = 5/325 per tablet.

   Number of tablets required = Ordered dose/Available dose

                                               = 10/650 / 5/325

                                               = 2 tablets q6h.